OpenStudy (anonymous):

2x^(2)-x-3=0 i need work shown step by step please

OpenStudy (anonymous):

The constant is negative you the signs of the factors will be + and -. What numbers multiply to get -3. (3, -1) or (-3, 1). Now, we will look to see which set of numbers can we multiply one by 2 (b/c of the 2x^(2)) and add the other numbers and get -1. The looks like -3 + 2(1) = 1. So we have our factors now. (2x-3)(x+1)=0 Set each one = 0 2x-3=0 x= 3/2 x+1=0 x= -1 Answer is x = 3/2, -1 You can input into original equation to double check.